OpenStudy (anonymous):

2. Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? Examples and a walkthrough please :)

OpenStudy (anonymous):

if you know one angle and the 90 degree angle, then you know two angles therefore you know the third angle (since they add up to 180)

OpenStudy (anonymous):

therefore in total you know all angles and one side, you can find the other two sides using trigonometry

OpenStudy (anonymous):



OpenStudy (anonymous):

in the triangle above, knowing that one leg is 7 we can get the other leg by putting \(\frac{7}{x}=\tan(20)\) and so \(x=\frac{7}{\tan(20)}\)

OpenStudy (ganpat):

 Sin x = b/h sin (90 - x) = a/h.. and so u can manipulate the formulas.. :)